Stage 1 ABC at a College: Assigning Costs to Activities
An accounting professor at Middleton University devotes 60 percent of her time to teaching, 30 percent of her time to research and writing, and 10 percent of her time to service activities such as committee work and student advising. The professor teaches two semesters per year. During each semester, she teaches one section of an introductory financial accounting course (with a maximum enrollment of 40 students) and one section of a graduate financial accounting course (with a maximum enrollment of 25 students). Including course preparation, classroom instruction, and appointments with students, each course requires an equal amount of time. The accounting professor is paid $120,000 per year.

Determine the activity cost of instruction per student in both the introductory and the graduate financial accounting courses.

Round answers to two decimal places.
Answer #1

Solution:-

Activity cost of instruction per student in the introductory financial accounting courses:-

$ 450 per student

Explanation:-

(120,000 * 0.60) / 2 = 36,000

= 36,000 / 2 = 18,000

= 18,000 / 40 = 450

Activity cost of instruction per student in the graduate financial accounting courses:-

$ 720 per student

Explanation:-

(120,000 * 0.60) / 2 = 36,000

= 36,000 / 2 = 18,000

= 18,000 / 25 = 720